Ike sells grapes at a farmers market. Each pound of grapes sells for the same amount. The table below shows the total prices for different numbers of pounds of grapes sold.
Number of Pounds Sold
2
4
10
Total
Price ($)

$4.50
$9.00
$22.50
A. In equation form describe the relationship between the number of pounds of grapes sold(x) and the total price (y), in dollars.

Answer by Boreal(15235)

Set it up as an ordered pair
(2, 4.50)
(4, 9)
(10, 22.50)
The slope is 4.50/2 or 2.25
slope-intercept form is y-y1=m(x-x1), m slope, (x1, y1) a point
so y-9=2.25(x-4)
y=2.25x
no constant.
each pound of grape is $2.25.
-
Can also be done by inspection. 10 pounds-22.50 is $2.25 per pound.
2 pounds for $4.50 is the same.